X y D z - Write the boolean expression:
Answer #1

Ans)

The above circuit digram is a combination of OR gate,NOT gate and AND gate,the X and Y inputs are passed through the OR gate which produces the result X+Y,the Z input is inverted by using the NOT gate and both thes are provided to the AND gate finally,hence the output Y= (X+Y).(Z'),consider the image for the clear solution with proper diagramsGiven circuit diagram, X+Y (x+y?) Z- 2 solution :- OR cate X x+y Y. NOT cate z 1 (X+Y) (2) Thug) y = (x+y) • ( ² )Please upvote if you liked my answer